# Self is the Best Learner: CT-free Ultra-Low-Dose PET Organ Segmentation via Collaborating Denoising and Segmentation Learning

### v1

@@ Abstract
Organ segmentation in Positron Emission Tomography (PET) plays a vital role in cancer quantification. Low-dose PET (LDPET) provides a safer alternative by reducing radiation exposure. However, the inherent noise and blurred boundaries make organ segmentation more challenging. Additionally, existing PET organ segmentation methods rely on co-registered Computed Tomography (CT) annotations, overlooking the problem of modality mismatch. In this study, we propose LDOS, a novel CT-free ultra-LDPET organ segmentation pipeline. Inspired by Masked Autoencoders (MAE), we reinterpret LDPET as a naturally masked version of Full-Dose PET (FDPET). LDOS adopts a simple yet effective architecture: a shared encoder extracts generalized features, while task-specific decoders independently refine outputs for denoising and segmentation. By integrating CT-derived organ annotations into the denoising process, LDOS improves anatomical boundary recognition and alleviates the PET/CT misalignments. Experiments demonstrate that LDOS achieves state-of-the-art performance with mean Dice scores of 73.11% (18F-FDG) and 73.97% (68Ga-FAPI) across 18 organs in 5% dose PET. Our code is publicly available.
@@ 1 Introduction
Positron emission tomography (PET) is a powerful molecular imaging that visualizes radiotracer distribution to reveal physiological processes. Low-dose PET (LDPET) reduces radiation exposure and has demonstrated diagnostic equivalence to Full-Dose PET (FDPET) [25, 26, 14, 4]. However, its clinical adoption remains limited by the lack of robust quantitative tools in high-noise settings. Accurate tracer uptake measurements in non-tumoral organs are critical for tumor quantification within Volumes of Interest (VOIs) [1, 15]. LDPET organ segmentation offers a promising approach for tracer uptake assessment and kinetic measurements [23, 27], yet it remains relatively understudied.
@@ 3.1.1 Shared Encoder
LDOS uses the nnU-Net backbone [9] with the ResEncL configuration [10]. Inspired by MAE, LDOS employs the LDPET denoising process to learn semantic latent representations. MAE uses the patch-level discretized images as inputs [7]. In the medical imaging context, large-scale patch masking leads to excessive information loss. In our work, LDOS is regarded as a pixel-level masking approach to retain finer details:
@@
[…]
@@ 3.1.2 Segmentation and Denoising Decoders
The encoder output feeds into both segmentation and denoising decoders, which share a consistent architecture. To enhance feature learning, a deep supervision strategy is incorporated:
$P_{full-dose}^{z}=D(\vartheta_{d}^{z};F({{x}_{i}}))$ (3)
$P_{seg}^{z}=\frac{\exp({{D}_{c}}(\vartheta_{s}^{z};F({{x}_{i}})))}{\sum\limits_{C}{\exp({{D}_{c}}(\vartheta_{s}^{z};F({{x}_{i}}))}}$ (4)
where $P_{full-dose}^{z}$ is the FDPET prediction at scale $z$, $D$ is the decoder, and $\vartheta_{d}^{z}$ denotes the decoder parameters. Similarly, $P_{seg}^{z}$ refers to the segmentation prediction for the $c$-th class, with ${{D}_{c}}$ being the decoder for class $c$ and $C$ the total number of segmentation classes.
@@ 4.1 Datasets and Implementation Details
LDOS was evaluated on datasets of images from patients injected with 18F-FDG and 68Ga-FAPI tracers, with sampling times of 300 seconds (100% dose) for FDPET and 15 seconds (5% dose) for ultra-LDPET. Each dataset included 18 segmented organs and was split into training (80%), validation (10%), and testing (10%) subsets.
LDOS was implemented based on the nnU-Net and trained from scratch. The patch size was $192\times 192\times 192$, and the batch size was 2. Training was conducted on an NVIDIA GeForce RTX 4090 GPU with 24 GB memory.
@@ 4.2 Multiorgan Segmentation Results
LDOS was compared to the segmentation results of real LDPET and FDPET. For a fair comparison, the same nnU-Net architecture was employed, excluding the denoising decoder loss. As shown in  2 and Fig. 4, LDOS achieved notable improvements in Dice scores, with $5.08\%$ and $8.24\%$ higher accuracy than LDPET on the 18F-FDG and 68Ga-FAPI datasets, respectively. Moreover, LDOS achieved segmentation accuracy comparable to or surpassing FDPET for certain organs.
@@
PET organ annotations for model training are derived from corresponding CT images, as direct organ labeling on PET images is impractical. However, achieving precise alignment of training is challenging due to inherent spatial shifts. As illustrated in Fig. 3, LDOS alleviates the misalignments by semantic feature learning of the self-denoising process, providing a novel solution to improve ultra-LDPET organ segmentation.
Table 4: Ablation studies of LDOS. LDPET-LDOS/ow means LDOS without the denoising process, and Fake-FDPET means that denoising and segmentation are performed sequentially.The best performance is highlighted in bold.
@@ 4.4 Ablation Experiment
To evaluate the contribution of the proposed pipeline, we compared its against LDPET and fake FDPET. As shown in Table 4, LDOS improved Dice scores on both the 18F-FDG and 68Ga-FAPI datasets. Unlike fake FDPET, which employed sequential denoising and segmentation, LDOS’s collaborative design achieved superior results without complex denoising model training. LDOS reduces deployment costs while enhancing usability. Additionally, results from the LDPET-LDOS/ow LDPET-LDOS/ow (omits the L1 loss weight attenuation strategy) indicates that focusing on denoising semantic information during the early training stages and prioritizing segmentation optimization in the later stages positively impacts the model’s overall performance.

### v2

@@ Abstract
Organ segmentation in Positron Emission Tomography (PET) plays a vital role in cancer quantification. Low-dose PET (LDPET) provides a safer alternative by reducing radiation exposure. However, the inherent noise and blurred boundaries make organ segmentation more challenging. Additionally, existing PET organ segmentation methods rely on co-registered Computed Tomography (CT) annotations, overlooking the problem of modality mismatch. In this study, we propose LDOS, a novel CT-free ultra-LDPET organ segmentation pipeline. Inspired by Masked Autoencoders (MAE), we reinterpret LDPET as a naturally masked version of Full-Dose PET (FDPET). LDOS adopts a simple yet effective architecture: a shared encoder extracts generalized features, while task-specific decoders independently refine outputs for denoising and segmentation. By integrating CT-derived organ annotations into the denoising process, LDOS improves anatomical boundary recognition and alleviates the PET/CT misalignments. Experiments demonstrate that LDOS achieves state-of-the-art performance with mean Dice scores of 73.11% (18F-FDG) and 73.97% (68Ga-FAPI) across 18 organs in 5% dose PET. Our code will be available at https://github.com/yezanting/LDOS.
@@ 1 Introduction
Positron emission tomography (PET) is a powerful molecular imaging modality that visualizes radiotracer distribution to reveal physiological processes. Low-dose PET (LDPET) reduces radiation exposure and has demonstrated diagnostic equivalence to Full-Dose PET (FDPET) [27, 28, 16, 5]. However, its clinical adoption remains limited by the lack of robust quantitative tools in high-noise settings. Accurate tracer uptake measurements in non-tumoral organs are critical for tumor quantification within Volumes of Interest (VOIs) [1, 17]. LDPET organ segmentation offers a promising approach for tracer uptake assessment and kinetic measurements [25, 29], yet it remains relatively understudied.
@@ 3.1.1 Shared Encoder
LDOS uses the nnU-Net backbone [11] with the ResEncL configuration [12]. Inspired by MAE, LDOS employs the LDPET denoising process to learn semantic latent representations. MAE uses the patch-level discretized images as inputs [8]. In the medical imaging context, large-scale patch masking leads to excessive information loss. The degradation in LDPET physically stems from lower photon counts and consequent statistical noise. LDOS approximates this effect as a pixel-level masking process, which facilitates the retention of core information:
@@
where ${{x}_{i}}$ is the LDPET image, ${{r}_{i}}$ is the FDPET image, and $\Lambda r$ represents the pixel-level discretization. The encoder processes the input ${{x}_{i}}$ to extract semantic latent features:
@@ 3.1.2 Segmentation and Denoising Decoders
The encoder output feeds into both segmentation and denoising decoders, which share a consistent architecture. To enhance feature learning, a deep supervision strategy is incorporated  [11]:
$P_{full-dose}^{z}=D_{d}(\vartheta_{d}^{z};F({{x}_{i}}))$ (3)
$P_{seg}^{z}=\frac{\exp({{D}_{s}^{c}}(\vartheta_{s}^{z};F({{x}_{i}})))}{\sum\limits_{C}{\exp({{D}_{s}^{c}}(\vartheta_{s}^{z};F({{x}_{i}}))}}$ (4)
where $P_{full-dose}^{z}$ is the FDPET prediction at scale $z$, $D$ is the decoder, and $\vartheta_{d}^{z}$ denotes the decoder parameters. Similarly, $P_{seg}^{z}$ refers to the segmentation prediction for the $c$-th class, with ${{D}_{s}^{c}}$ being the decoder for class $c$ and $C$ the total number of segmentation classes.
@@ 4.1 Datasets and Implementation Details
Excluding significant misalignments, we used 52 18F-FDG and 60 68Ga-FAPI tracer scans acquired from Nanfang Hospital Southern Medical University on a UIH uEXPLORER (Total-Body) PET/CT to validate LDOS. The sampling times were set to 300 seconds (100% dose) for FDPET and 15 seconds (5% dose) for ultra-LDPET. Each dataset included 18 segmented organs and was split into training (80%), validation (10%), and testing (10%) subsets. Five-fold cross-validation was performed separately per tracer. Notably, direct organ annotation on PET images was challenging due to constraints in the clinical workflow. Consequently, we utilized annotations from corresponding CT scans as the gold standard, which were then manually refined. Once trained, our model performs organ segmentation on PET images alone, without reliance on the CT modality.
LDOS was implemented based on the nnU-Net and trained from scratch. We trained 500 epoches. The patch size was $192\times 192\times 192$, and the batch size was 2. Training was conducted on an NVIDIA GeForce RTX 4090 GPU with 24 GB memory.
@@ 4.2 Multiorgan Segmentation Results
LDOS was compared to the segmentation results of real LDPET and FDPET. For a fair comparison, the same nnU-Net architecture was employed, excluding the denoising decoder loss. As shown in Table 2 and Fig. 4, LDOS achieved notable improvements in Dice scores, with $5.08\%$ and $8.24\%$ higher accuracy than LDPET on the 18F-FDG and 68Ga-FAPI datasets, respectively. Moreover, LDOS achieved segmentation accuracy comparable to or surpassing FDPET for certain organs.
@@
PET organ annotations for model training are derived from corresponding CT images, as direct organ labeling on PET images is impractical. However, achieving precise alignment of training is challenging due to inherent spatial shifts. As illustrated in Fig. 3, LDOS mitigates the misalignments by semantic feature learning of the self-denoising process, providing a novel solution to improve ultra-LDPET organ segmentation.
Table 4: Ablation studies of LDOS. LDPET-LDOS/ow means LDOS without the L1 loss weight attenuation strategy, and Fake-FDPET means that denoising and segmentation are performed sequentially. The best performance is highlighted in bold.
@@ 4.4 Ablation Experiment
To evaluate the contribution of the proposed pipeline, we compared its against LDPET and fake FDPET. As shown in Table 4, LDOS improved Dice scores on both the 18F-FDG and 68Ga-FAPI datasets. Unlike fake FDPET, which employed sequential denoising and segmentation, LDOS’s collaborative design achieved superior results without complex denoising model training. LDOS reduces deployment costs while enhancing usability. Additionally, results from the LDPET-LDOS/ow (omits the L1 loss weight attenuation strategy) indicates that focusing on denoising semantic information during the early training stages and prioritizing segmentation optimization in the later stages positively impacts the model’s overall performance.
@@ 5 Conclusion
{credits}
5.0.1 Acknowledgements
This research was funded by the National Natural Science Foundation of China (62371221, 12326616 and 62201245), National High end Foreign Experts Recruitment Plan (G2023030025L)and the Science and Technology Program of Guangdong Province 2022A0505050039.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.